OpenStudy (anonymous):

If the y in y=mx+b is more than 1, such as 2y=mx+b or 3y=mx+b what do you do.

OpenStudy (anonymous):

Is it possible to divide both sides? Because I was taught that u cant divide an x by a y.

OpenStudy (anonymous):

what was the equation of the line in point slope form (i.e. y-y1=m(x-x1)?

OpenStudy (anonymous):

It says im supposed to graph 3y=4x-5 and I just dont know how to get rid of the 3 in the 3y

OpenStudy (anonymous):

If you just need to graph, use this website(: Type in 3y=4x-5 http://www.mathsisfun.com/data/grapher-equation.html

OpenStudy (anonymous):

well, logically the 4x is the slope, and the -5 is the y-int, so in order to get rid of the 3 you would divide both the 4 and the -5 by 3. That will keep it in y=mx+b form

OpenStudy (anonymous):

so y=4/3x-5/3 is the equation for the line.

OpenStudy (anonymous):

Thanks dude It all makes sense now

OpenStudy (anonymous):

yw